# compat=1.0.0

### v1

@@ 4.2 $n-\overline{n}$  Oscillation
$\mathcal{L}_{Y}\supset Y_{6i}\;\overline{5}_{Fi}24_{F}45_{H}\supset\overline{Y}_{i}\;d^{c}_{i}\xi_{3}\Sigma_{6},$ (4.8)
where $\overline{Y}_{i}=\frac{1}{2}\left(D^{T}_{c}Y_{6}^{T}\right)_{i}$ is written in the physical basis. Of course, it is only the value of $\overline{Y}_{1}$ that matters for our purposes. The upper two vertices in Fig. 8, on the other hand, are defined in Eq. (3.8).
[…]
@@
where the relevant transition matrix element typically has the value $\mathcal{M}\sim 10^{-5}\,\mathrm{GeV}^{6}$. (See, for example, Table 1 of Ref. [46] for more details.) If we assume $m_{\xi_{3}}=m_{\Sigma_{6}}=M$ and set all relevant Yukawa couplings to one, Eq. (4.10) yields $\tau_{n-\overline{n}}\gtrsim 4.7\times 10^{8}\,\mathrm{sec}$ for $M\gtrsim 3.7\times 10^{5}\,\mathrm{GeV}$. Finally, if we demand an observability of the $n-\overline{n}$ oscillation signal in near future, the relevant range for Model B reads
@@ 4.3.2 Numerical fit
$\sum_{i}m_{i}$ (meV) <(8787120120[74] 59 59
$m_{\beta\beta}$ (meV) <(2828122122[75] 3.69 48.4
$\delta^{\mathrm{PMNS}}_{\mathrm{CP}}$ (deg) $177^{+19}_{-20}$$285^{+25}_{-28}$ [76] 191.67 174.039
Table 3: Charged fermion mass parameter input at $M_{\mathrm{GUT}}$ is taken from Ref.  [73]. Neutrino observables are taken from Refs. [77, 76]. The Model B (NO) fit yields $\chi^{2}=1.0$, whereas Model B (IO) fit gives $\chi^{2}=0.13$.
Predictions for both benchmark scenarios within Model B framework are presented in Fig. 9. The pink shaded part in Fig. 9 highlights the current LHC bound due to dijet resonance. Vertical yellow shaded region is a constraint due to pair-produced octet colored fermions, where each fermion decays into three jets, thus mimicking R-parity violating gluino signals [79].
Figure 9: Observable $n-\overline{n}$ ranges for the Model B (NO) and Model B (IO) scenarios. The shaded blue and magenta regions correspond to the viable parts of the parameter space that can be probed in the future by NNBAR experiment and DUNE, respectively, whereas the shaded gray region plotted on top of magenta and blue regions shows the current exclusion from SK. The pink shaded part highlights the current LHC bound due to dijet resonance, whereas the vertical yellow shaded region corresponds to exclusion from pair-produced octet colored fermions, each decaying into three jets [79].
@@
  • These scenarios, specified by the benchmark fits presented above, can be tested in upcoming neutrino experiments through their predictions for neutrino mixing parameters, notably at DUNE [80], JUNO [81], and Hyper-Kamiokande [82]. In addition, future searches for neutrinoless double-beta decay, such as nEXO [83], will fully probe the relevant parameter space in both models. Moreover, the sum of neutrino masses is already stringently constrained by cosmological measurements [74], and, in the near future, the inverted mass ordering scenario for Model B may be fully ruled out.
  • For the models presented in this work, future searches for $n-\overline{n}$ oscillations at DUNE and at the NNBAR facility at ESS are expected to achieve unprecedented sensitivities, probing new-physics scales up to $\sim 10^{10}$ GeV and $\sim 10^{11}$ GeV, respectively.
@@ 5 Conclusions
Neutron-antineutron oscillations offer a powerful probe of baryon number violation and high-scale new physics well beyond the Standard Model. Future experiments such as DUNE and the proposed NNBAR at ESS are poised to extend sensitivity to such $\Delta B=2$ processes at unprecedented levels, probing energy scales far beyond the reach of colliders. Motivated by this, we have presented two simple $SU(5)$-based models in which the embedding of the seesaw mechanism for neutrino masses naturally realizes two distinguishable topologies for $n-\overline{n}$ oscillations. One is mediated by color-sextet scalars (arising in a Type II seesaw context), and the other one involves both a scalar sextet and a color-octet fermion (emerging from a Type III seesaw structure). While the first topology can also arise within the SO(10)SO(10)/Pati–Salam framework, the second embeds naturally in $SU(5)$, this distinction constitutes a key novelty of our proposal. In both models, the same dynamics responsible for generating fermion masses also induces baryon number violation, linking flavor structure and baryon number violation in a unified framework. These findings highlight the unique potential of $n-\overline{n}$ oscillation searches to access ultra-heavy new physics and offer a rare low-energy portal to grand unification. We also demonstrated that a consistent gauge coupling unification, in agreement with current proton decay bounds, is achieved within these models in the parameter space where $n-\overline{n}$ oscillations are potentially observable. Gauge-mediated proton decay, if observed, would offer a complementary probe for testing these models.
@@ Acknowledgments
I.D. acknowledges the financial support from the Slovenian Research Agency (research core funding No. P1-0035). S.F. and S.S. also acknowledge the financial support from the Slovenian Research Agency (research core funding No. P1-0035 and N1-0321).

### v2

@@ 4.2 $n-\overline{n}$  Oscillation
$\mathcal{L}_{Y}\supset Y_{6i}\;\overline{5}_{Fi}24_{F}45_{H}\supset\overline{Y}_{i}\;d^{c}_{i}\xi_{3}\Sigma_{6}\;\varepsilon_{c},$ (4.8)
where $\overline{Y}_{i}=\frac{1}{2}\left(D^{T}_{c}Y_{6}^{T}\right)_{i}$ is written in the physical basis ($\varepsilon_{c}$ is the three-index Levi-Civita tensor). The gauge invariance of the term on the right-hand side can be understood from the fact that, under the color $SU(3)_{c}$ group, $\overline{6}\times\overline{3}\supset 8$. Of course, it is only the value of $\overline{Y}_{1}$ that matters for our purposes. The upper two vertices in Fig. 8, on the other hand, are defined in Eq. (3.8).
$d^{c}$$u^{c}$$u^{c}$$d^{c}$$d^{c}$$d^{c}$$\Sigma^{*}_{6}$$\Sigma^{*}_{6}$$\xi_{3}$$\overline{Y}_{1}$$\overline{Y}_{1}$$Y^{ud}_{11}$$Y^{ud}_{11}$
@@
where the relevant transition matrix element typically has the value $\mathcal{M}\sim 10^{-5}\,\mathrm{GeV}^{6}$. (See, for example, Table 1 of Ref. Fridell:2021gag for more details.) If we assume $m_{\xi_{3}}=m_{\Sigma_{6}}=M$ and set all relevant Yukawa couplings to one, Eq. (4.10) yields $\tau_{n-\overline{n}}\gtrsim 4.7\times 10^{8}\,\mathrm{sec}$ for $M\gtrsim 3.7\times 10^{5}\,\mathrm{GeV}$. Finally, if we demand an observability of the $n-\overline{n}$ oscillation signal in near future, the relevant range for Model B reads
@@ 4.3.2 Numerical fit
$\sum_{i}m_{i}$ (meV) <(8787120120ParticleDataGroup:2024cfk 59 59
$m_{\beta\beta}$ (meV) <(2828122122KamLAND-Zen:2024eml 3.69 48.4
$\delta^{\mathrm{PMNS}}_{\mathrm{CP}}$ (deg) $177^{+19}_{-20}$$285^{+25}_{-28}$ NUFIT 191.67 174.039
Table 3: Charged fermion mass parameter input at $M_{\mathrm{GUT}}$ is taken from Ref. Babu:2016bmy . Neutrino observables are taken from Refs. Esteban:2024eli , NUFIT . The Model B (NO) fit yields $\chi^{2}=1.0$, whereas Model B (IO) fit gives $\chi^{2}=0.13$.
Predictions for both benchmark scenarios within Model B framework are presented in Fig. 9. The pink shaded part in Fig. 9 highlights the current LHC bound due to dijet resonance. Vertical yellow shaded region is a constraint due to pair-produced octet colored fermions, where each fermion decays into three jets, thus mimicking R-parity violating gluino signals ATLAS:2024kqk .
Figure 9: Observable $n-\overline{n}$ ranges for the Model B (NO) and Model B (IO) scenarios. The shaded blue and magenta regions correspond to the viable parts of the parameter space that can be probed in the future by NNBAR experiment and DUNE, respectively, whereas the shaded gray region plotted on top of magenta and blue regions shows the current exclusion from SK. The pink shaded part highlights the current LHC bound due to dijet resonance, whereas the vertical yellow shaded region corresponds to exclusion from pair-produced octet colored fermions, each decaying into three jets ATLAS:2024kqk . Projected sensitivity for dijet resonances from FCC-$h$ is presented by the dash-dotted red line, which will probe a significant part of the existing parameter space.
@@
  • These scenarios, specified by the benchmark fits presented above, can be tested in upcoming neutrino experiments through their predictions for neutrino mixing parameters, notably at DUNE DUNE:2020jqi , JUNO JUNO:2015zny , and Hyper-Kamiokande Hyper-Kamiokande:2018ofw . In addition, future searches for neutrinoless double-beta decay, such as nEXO nEXO:2021ujk , will fully probe the relevant parameter space in both models. Moreover, the sum of neutrino masses is already stringently constrained by cosmological measurements ParticleDataGroup:2024cfk , and, in the near future, the inverted mass ordering scenario for Model B may be fully ruled out.
  • For the models presented in this work, future searches for $n-\overline{n}$ oscillations at DUNE and at the NNBAR facility at ESS are expected to achieve unprecedented sensitivities, probing new-physics scales up to $10^{10}$ GeV and $10^{11}$ GeV, respectively, while the other state involved in the process needs to be near the TeV scale.
  • It should be understood that the non-observation of $n-\overline{n}$ oscillation would not rule out the model. It would only indicate that the states responsible for this $\Delta B=2$ process are relatively heavy, leading to a suppression of the $n-\overline{n}$ transition rate.
@@ 5 Conclusions
Neutron-antineutron oscillations offer a powerful probe of baryon number violation and high-scale new physics well beyond the Standard Model. Future experiments such as DUNE and the proposed NNBAR at ESS are poised to extend sensitivity to such $\Delta B=2$ processes at unprecedented levels, probing energy scales far beyond the reach of colliders. Motivated by this, we have presented two simple $SU(5)$-based models in which the embedding of the seesaw mechanism for neutrino masses naturally realizes two distinguishable topologies for $n-\overline{n}$ oscillations. One is mediated by color-sextet scalars (arising in a Type II seesaw context), and the other one involves both a scalar sextet and a color-octet fermion (emerging from a Type III seesaw structure). While the first topology can also arise within the SO(10)SO(10)/Pati–Salam framework, the second embeds naturally in $SU(5)$, this distinction constitutes a key novelty of our proposal. In both models, the same dynamics responsible for generating fermion masses also induces baryon number violation, linking flavor structure and baryon number violation in a unified framework. Moreover, we have shown that if $n-\overline{n}$ oscillations are observed, the FCC-hh’s projected sensitivity to dijet resonances will probe a substantial portion of the allowed parameter space. These findings highlight the unique potential of $n-\overline{n}$ oscillation searches to access ultra-heavy new physics and offer a rare low-energy portal to grand unification. We also demonstrated that a consistent gauge coupling unification, in agreement with current proton decay bounds, is achieved within these models in the parameter space where $n-\overline{n}$ oscillations are potentially observable. Gauge-mediated proton decay, if observed, would offer a complementary probe for testing these models.
@@ Acknowledgments
I.D. acknowledges the financial support from the Slovenian Research Agency (research core funding No. P1-0035). S.F. and S.S. acknowledge the financial support from the Slovenian Research Agency (research core funding No. P1-0035 and N1-0321).